# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

### v1

@@ Abstract
The generative adversarial network (GAN) aims to approximate an unknown distribution via a parameterized neural network (NN). While GANs have been widely applied in reinforcement and semisupervised learning as well as computer vision tasks, selecting their parameters often needs an exhaustive search and only a few selection methods can be proved to be theoretically optimal. One of the most promising GAN variants is the Wasserstein GAN (WGAN). Prior work on optimal parameters for WGAN is limited to the linear-quadratic-Gaussian (LQG) setting, where the NN is linear and the data is Gaussian. In this paper, we focus on the characterization of optimal WGAN parameters beyond the LQG setting. We derive closed-form optimal parameters for one-dimensional WGANs when the NN has non-linear activation functions and the data is non-Gaussian. To extend this to high-dimensional WGANs, we adopt the sliced Wasserstein framework and replace the constraint on marginal distributions of the randomly projected data by a constraint on the joint distribution of the original (unprojected) data. We show that the linear generator can be asymptotically optimal for sliced WGAN with non-Gaussian data. Empirical studies show that our closed-form WGAN parameters have good convergence behavior with data under both Gaussian and Laplace distributions. Also, compared to the r principal component analysis (r-PCA) solution, our proposed solution for sliced WGAN can achieve the same performance while requiring less computational resources.
@@ I Introduction
Generative adversarial networks (GANs) are machine learning frameworks put forth by Goodfellow et al. [1]. A GAN aims to learn an unknown distribution from training data via two competing components, namely the generator and the discriminator. The former tries to mimic the distribution of training data while the latter discriminates between true data and generated data. This framework formulates a minimax problem, where the discriminator aims to maximize the probability of differentiating true data from generated ones, while the generator seeks to minimize this probability [1]. Besides computer vision tasks, applications of GANs to communication systems have also received a lot of attentions. For example, GANs have been applied to autonomous wireless channel modeling [2][3] and mmWave MIMO channel estimation [4].
@@
Traditionally, both the generator and discriminator in GAN are approximated by neural networks (NNs) [1, 5]. Without the NN restrictions and under an optimal unconstrained discriminator, the minimax problem associated with a GAN becomes the minimization of the Jensen-Shannon divergence (JSD) between the distributions of true and generated data. However, due to the nature of this minimax game, the GAN suffers from several problems including vanishing gradient and mode collapse. Many variants of GAN have been proposed to solve these problems, and one of the most promising variants is the Wasserstein GAN (WGAN) [6] , which replaces the JSD by the Wasserstein distance from the optimal transport theory [7]. The WGAN is differentiable with respect to the generator parameters almost everywhere, which benefits the convergence of stochastic gradient descent (SGD) widely adopted for training NNs [6].
@@
Despite the substantial success of applying WGAN to learn distributions in real applications, there are only a few GAN parameter selection algorithms proved to be theoretically optimal [8, 9], which limits the development of GAN beyond heuristic methods in [1, 5]. This lack of rigorous analysis also restricts the evaluation of GANs’ performance to subjective terms. One exception is [8] where Feizi et al. attempted to theoretically understand WGANs on a simple linear quadratic Gaussian (LQG) setting. In this benchmark setting, the synthetic data is generated by a Gaussian distribution, the generator NN is restricted to be linear, and the loss function is quadratic. It is shown in [8, Theorem 1] that for this simple setting, the optimal GAN solution happens to be the principal component analysis (PCA) solution. Regularized versions of GANs are also considered and analyzed [5]. Optimal WGAN solutions under LQG settings, with additional entropic and Sinkhorn regularizers, are also studied [9].
@@
In this paper, we aim to analytically solve WGANs beyond the LQG setting. As described in Sec. II, our setting allows non-Gaussian data distribution and non-linear generators constructed from sigmoid and ReLU functions, and is therefore more general than [8, 9, 10]. Also, we attempt to exactly solve WGANs rather than their regularized versions as in [9]. All of these make our problem much more challenging—after all, even for the inner discriminator problem, which is an optimal transport problem, the solution in most cases is numerically approximated but not analytically characterized [7]. To overcome the challenge, we first focus on one-dimensional data and generator in Sec. III-A, where we provide closed-form solutions for optimal generators in one-dimensional WGANs defined in Sec. II. For high-dimensional WGANs, we adopt the sliced WGAN framework and define a revised sliced WGAN with unprojected data distribution in Sec. II. We provide closed-form solutions for asymptotically optimal generators in high-dimensional sliced WGAN in III-B. The proofs are presented in Sec. IV, where we leverage on results in [11] to solve the inner discriminator problem in closed-form, which greatly simplifies the necessary conditions of optimal WGAN parameters. Moreover, our closed-form solutions do not need any training for the discriminator as [12] and hence provide additional benefit for training WGAN with a decentralized system [13]. Empirical studies in Sec. V show that our closed-form one-dimensional WGAN parameters have good convergence behavior with synthetic data under both Gaussian and Laplace distributions. Empirical studies also show that our closed-form high-dimesional sliced WGAN can acheive the same performance as r-PCA solutions, while requiring fewer computational resources.
@@
Throughout the paper, we adopt the following notational conventions. The operator $\mathrm{diag}(\cdot):\mathbb{R}^{d\times d}\to\mathbb{R}^{d}$ extracts the diagonal entries of a $d\times d$ matrix and returns them as a $d$-dimensional vector. We let $\mathbf{1}\in\mathbb{R}^{d}$ denote the vector whose components are all equal to one, and $e_{i}\in\mathbb{R}^{d}$ denote the $i$-th standard unit vector, with its $i$-th component equal to $1$ and all others equal to 0.
@@ II Problem Formulation
For the WGAN considered in this paper, the generator function is denoted by $G_{\Theta}(.)$, where $\Theta$ is the generator NN’s parameter (weights). Let $q\in\{1,2\}$ represent the order of the Wasserstein distance. In a $q^{\mathrm{th}}$-order WGAN setting, one aims to solve
@@
for an optimal parameter $\Theta$, where $\|\cdot\|$ denotes the $\ell_{2}$-norm in $\mathbb{R}^{d}$, $\mu$ and $\nu^{\Theta}$ are probability measures on $\mathbb{R}^{d}$, generated by the data $X$ and the generator output $G_{\Theta}(Z)$ for a given $\Theta$ and Gaussian input $Z$, respectively. Also, $\Pi(\mu,\nu^{\Theta})$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginal constraints $\mu$ and $\nu^{\Theta}$, respectively, satisfying $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{q}d\mu(x)d\nu^{\Theta}(y)<\infty.$ The WGAN problem described in (1) can be equivalently written as $\min_{\Theta}(\mathbb{E}_{\mu}[P(X,\Theta)])^{1/q}$ where the inner-discriminator problem is defined as
@@
Note that (2) belongs to the family of the optimal transport problems with $q^{\mathrm{th}}$-order Wasserstein distance, $q\in\{1,2\}$ [7, Proposition 2.2]. In this paper, we consider the following activation functions: 1) the linear function; 2) the rectified linear unit (ReLU) function $\max(0;z)$; and 3) the sigmoid function $1/(1+\exp(-z)).$
Following [1][8], (1) is a population WGAN problem, where $\Theta$ can be optimized with the true data distribution $\mu$, and studying the theoretical population WGAN can lead to practical design with large enough empirical data. In the next Sec. III-A, we first theoretically solve the population WGAN and obtain the closed-form solutions under $d=1$.
Note that for WGAN (1), the closed-form solution was only obtained in the LQG setting [8]. In the LQG setting, the $d$-dimensional synthetic data is generated by Gaussian distribution, i.e., $X\sim\mathcal{N}(0,\mathbf{K})$, the generator is restricted to be a linear generator of the form
@@
Beyond the LQG setting, we show how to apply our $d=1$ results for (1) to cope with higher dimensional data $d>1$ by adopting the sliced Wasserstein distance, which offers a computationally efficient alternative to the WGAN [14, 15, 16]. Let $\Omega=\{\omega\in\mathbb{R}^{d}:~\|\omega\|=1\}$ contains all the directions in $\mathbb{R}^{d}$. In sliced WGAN, we project both the data and the generator’s output onto $\mathbb{R}$ through a random direction $\omega\in\Omega$ and compute the Wasserstein distance between the projected 1-dimensional distributions $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$, i.e., $\omega^{T}X$ and $\omega^{T}G_{\Theta}(Z)$ respectively follow distributions $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$. By replacing the Wasserstein distance of inner-discriminator problem (2) with sliced one,
$\mathbb{E}_{\mu_{\omega}}[P(X,\Theta)]:=\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega,$ (4)
@@
$\min_{\Theta}\mathbb{E}_{\mu_{\omega}}[P(X,\Theta)]^{\frac{1}{q}}$
$=\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\right)^{\frac{1}{q}}.$ (5)
Moreover, though widely adopted in [14, 15, 16], the inner-discriminator problem (4) for the sliced WGAN (II) is optimized under randomly projected data marginal constraint $\mu_{\omega}$. Compared with (2), it is reasonable to change the marginal constraint on data of $\pi$ from projected $\mu_{\omega}$ to unprojected $\mu$,
$\pi\in\Pi(\mu,\nu^{\Theta}_{\omega}),$
and the new sliced WGAN problem is
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\right)^{\frac{1}{q}}.$ (6)
Note that for original (II) and the new (6) with unprojected $\mu$, the optimal values are both non-negative. Thus, we say that a generator $G_{\Theta}(Z)$ is asymptotically optimal for (II) if the corresponding value goes to zero as $d\rightarrow\infty$. The same definition applies to (6). In Sec. III-B, we identify the asymptotically optimal generator for both original and unprojected sliced WGAN problems (II) and (6).

### v2

@@ Abstract
The generative adversarial network (GAN) aims to approximate an unknown distribution via a parameterized neural network (NN). While GANs have been widely applied in reinforcement and semi-supervised learning as well as computer vision tasks, selecting their parameters often needs an exhaustive search, and only a few selection methods have been proven to be theoretically optimal. One of the most promising GAN variants is the Wasserstein GAN (WGAN). Prior work on optimal parameters for population WGAN is limited to the linear-quadratic-Gaussian (LQG) setting, where the generator NN is linear, and the data is Gaussian. In this paper, we focus on the characterization of optimal solutions of population WGAN beyond the LQG setting. As a basic result, closed-form optimal parameters for one-dimensional WGAN are derived when the NN has non-linear activation functions, and the data is non-Gaussian. For high-dimensional data, we adopt the sliced Wasserstein framework and show that the linear generator can be asymptotically optimal. Moreover, the original sliced WGAN only constrains the projected data marginal instead of the whole one in classical WGAN, and thus, we propose another new unprojected sliced WGAN and identify its asymptotic optimality. Empirical studies show that compared to the celebrated r-principal component analysis (r-PCA) solution, which has cubic complexity to the data dimension, our generator for sliced WGAN can achieve better performance with only linear complexity.
@@ I Introduction
Generative adversarial networks (GANs) are machine learning frameworks put forth by Goodfellow et al. [14]. A GAN aims to learn an unknown distribution from training data via two competing components, namely the generator and the discriminator. The former tries to mimic the distribution of training data while the latter discriminates between true data and generated data. This framework formulates a minimax problem, where the discriminator aims to maximize the probability of differentiating true data from generated ones, while the generator seeks to minimize this probability [14]. Besides computer vision tasks, applications of GANs to communication systems have also garnered significant attention. For example, GANs have been applied to autonomous wireless channel modeling [37, 21] and mmWave MIMO channel estimation [11].
@@
Traditionally, both the generator and discriminator in GAN are approximated by neural networks (NNs) [14, 30]. Without the NN restrictions and under an optimal unconstrained discriminator, the minimax problem associated with a GAN becomes the minimization of the Jensen-Shannon divergence (JSD) between the distributions of true and generated data. However, due to the nature of this minimax game, the GAN suffers from several problems, including vanishing gradient and mode collapse. Many variants of GAN have been proposed to solve these problems, and one of the most promising variants is the Wasserstein GAN (WGAN) [2], which replaces the JSD by the Wasserstein distance from the optimal transport theory [27]. The WGAN is differentiable with respect to the generator parameters almost everywhere, which benefits the convergence of stochastic gradient descent (SGD) widely adopted for training NNs [2].
@@
Despite the substantial success of applying WGAN to learn distributions in real applications, there are only a few GAN parameter selection algorithms proven to be theoretically optimal [12, 29], which limits the development of GAN beyond heuristic methods in [14, 30]. This lack of rigorous analysis also restricts the evaluation of GANs’ performance to subjective terms. One exception is [12] where Feizi et al. attempted to theoretically understand WGANs in a simple linear quadratic Gaussian (LQG) setting. In this benchmark setting, the synthetic data is generated by a Gaussian distribution, the generator NN is restricted to be linear, and the loss function is quadratic. It is shown in [12, Theorem 1] that for this simple setting, the optimal GAN solution happens to be the principal component analysis (PCA) solution. Regularized versions of GANs are also considered and analyzed [30]. Optimal WGAN solutions under LQG settings, with additional entropic and Sinkhorn regularizers, are also studied [29].
@@
In this paper, we aim to solve WGANs beyond the LQG setting analytically. As described in Sec. II, our setting allows non-Gaussian data distribution and non-linear generators constructed from sigmoid and ReLU functions, and is therefore more general than [12, 29, 3]. Also, we attempt to exactly solve WGANs rather than their regularized versions as in [29]. All of these make our problem much more challenging—after all, even for the inner discriminator problem, which is an optimal transport problem, the solution in most cases is numerically approximated but not analytically characterized [27]. To overcome the challenge, we first focus on one-dimensional data and generator in Sec. III-A, where we provide closed-form solutions for optimal generators in one-dimensional WGANs defined in Sec. II.
For high-dimensional WGANs, we adopt the sliced WGAN framework, which provides a computationally efficient alternative by reducing high-dimensional distributions to one-dimensional projections and averaging Wasserstein distances of projected distributions [10, 20, 25, 35]. Also, we define a revised sliced WGAN with unprojected data distribution in Sec. II. We provide closed-form solutions for asymptotically optimal generators in high-dimensional sliced WGAN in III-B. The proofs are presented in Sec. IV, where we leverage results in [1] to solve the inner discriminator problem in closed form, which significantly simplifies the necessary conditions for optimal WGAN parameters. Moreover, our closed-form solutions do not need any training for the discriminator as [8] and hence provide additional benefit for training WGAN with a decentralized system [24]. Empirical studies in Sec. V show that our closed-form one-dimensional WGAN parameters have good convergence behavior with synthetic data under both Gaussian and Laplace distributions. High-dimensional empirical studies also show that our closed-form sliced WGAN generator can achieve similar or better sliced-Wasserstein distance performance as r-PCA solutions in [12][13], while reducing computational complexity from cubic to linear in data dimension. Compared with our previous work [16], the asymptotic optimality of linear generators in Theorems 2 and 3 is new, as is the proof of Theorem 1, which is not presented in [16].
@@
Throughout the paper, we adopt the following notational conventions. The operator $\mathrm{diag}(\cdot):\mathbb{R}^{d\times d}\to\mathbb{R}^{d}$ extracts the diagonal entries of a $d\times d$ matrix and returns them as a $d$-dimensional vector. We let $\mathbf{1}\in\mathbb{R}^{d}$ denote the vector whose components are all equal to one, and $e_{i}\in\mathbb{R}^{d}$ denote the $i$-th standard unit vector, with its $i$-th component equal to $1$ and all others equal to 0. If $\mu$ is a probability measure on $\mathcal{X}$, and $T$ is a mapping $\mathcal{X}\rightarrow\mathcal{Y}$, then $T\#\mu$ stands for the push-forward of $\mu$ by $T$.
@@ II Problem Formulation
For the WGAN considered in this paper, the generator function is denoted by $G_{\Theta}(\cdot)$, where $\Theta$ are the generator NN’s parameters (weights). Let $q\in\{1,2\}$ represent the order of the Wasserstein distance. In a $q^{\mathrm{th}}$-order WGAN setting, one aims to solve
@@
for an optimal parameter $\Theta$, where $\|\cdot\|$ denotes the $\ell_{2}$-norm in $\mathbb{R}^{d}$, $\mu$ and $\nu^{\Theta}$ are probability measures on $\mathbb{R}^{d}$, generated by the data $X$ and the generator output $G_{\Theta}(Z)$ for a given $\Theta$ and Gaussian input $Z$, respectively. Also, $\Pi(\mu,\nu^{\Theta})$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginal constraints $\mu$ and $\nu^{\Theta}$, respectively. That is, a $\pi\in\Pi$ satisfies the marginal conditions: $(\text{proj}^{1}_{\pi}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d})\#\pi=\mu$, and $(\text{proj}^{2}_{\pi}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d})\#\pi=\nu^{\Theta}$, where $\text{proj}^{1}_{\pi}$ denotes the projection onto the first coordinate of $\pi$, and $\text{proj}^{2}_{\pi}$ denotes the projection onto the second coordinate of $\pi$. The marginal constraints $\mu$ and $\nu^{\Theta}$ satisfy $\int_{\mathbb{R}^{d}}\|x\|^{q}d\mu(x)<\infty,$ and $\int_{\mathbb{R}^{d}}\|y\|^{q}d\nu^{\Theta}(y)<\infty$. The WGAN problem described in (1) can be equivalently written as $\min_{\Theta}(\mathbb{E}_{\mu}[P(X,\Theta)])^{1/q}$ where the inner-discriminator problem is defined as
@@
Note that (2) belongs to the family of optimal transport problems with $q^{\mathrm{th}}$-order Wasserstein distance, $q\in\{1,2\}$ [27, Proposition 2.2]. In this paper, we consider the following activation functions: 1) the linear function; 2) the rectified linear unit (ReLU) function $\max(0;z)$; and 3) the sigmoid function $1/(1+\exp(-z)).$
Following [14][12], (1) is a population WGAN problem, where $\Theta$ can be optimized with the true data distribution $\mu$, and studying the theoretical population WGAN can lead to practical design with large enough empirical data. In Sec. III-A, we first theoretically solve the population WGAN and obtain the closed-form solutions under $d=1$.
Note that for WGAN (1), the closed-form solution was only obtained in the LQG setting [12]. In the LQG setting, the $d$-dimensional synthetic data is generated by a Gaussian distribution, i.e., $X\sim\mathcal{N}(0,\mathbf{K})$, the generator is restricted to be a linear generator of the form
@@
Beyond the LQG setting, we show how to apply our $d=1$ results for (1) to cope with higher-dimensional data $d>1$ by adopting the sliced Wasserstein distance, which offers a computationally efficient alternative to the WGAN [10, 20, 25]. Consider the unit sphere
$\Omega=\{\omega\in\mathbb{R}^{d}:~\|\omega\|=1\},$ (4)
which contains all the directions in $\mathbb{R}^{d}$. In sliced WGAN, we project both the data and the generator’s output onto $\mathbb{R}$ through a random direction $\omega\in\Omega$ and compute the Wasserstein distance between the projected one-dimensional distributions $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$, i.e., $\omega^{T}X$ and $\omega^{T}G_{\Theta}(Z)$ respectively follow distributions $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$. By replacing the Wasserstein distance of the inner-discriminator problem (2) with the sliced one,
$\mathbb{E}_{\mu_{\omega}}[P(X,\Theta)]:=\int_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega,$ (5)
@@
$\min_{\Theta}\mathbb{E}_{\mu_{\omega}}[P(X,\Theta)]^{\frac{1}{q}}=\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\right)^{\frac{1}{q}},$ (6)
where $\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})$ is the set of probability measures on $\mathbb{R}\times\mathbb{R}$ with marginal constraints $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$, generated by the projected data $\omega^{T}X$ and the projected generator output $\omega^{T}G_{\Theta}(Z)$, respectively. That is, a $\pi_{\omega}\in\Pi_{\omega}$ satisfies the marginal conditions: $(\text{proj}^{1}_{\pi_{\omega}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R})\#\pi_{\omega}=\mu_{\omega}$, and $(\text{proj}^{2}_{\pi_{\omega}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R})\#\pi_{\omega}=\nu^{\Theta}_{\omega}$.
Moreover, though widely adopted in [10, 20, 25], the joint PDF set $\Pi_{\omega}$ of the inner-discriminator problem (5) for the sliced WGAN (6) is constrained by projected marginal $\mu_{\omega}$ on $\omega^{T}X$. However, since projections of distributions may not be one-to-one, the optimal $\pi_{\omega}$ may correspond to a data distribution $\tilde{\mu}$ on $\mathbb{R}^{d}$ that is different from the original $\mu$, while $\tilde{\mu}$ and $\mu$ have the same $\mu_{\omega}$ after projection. To avoid this issue, it is reasonable to consider the set of probability measures $\Pi^{\prime}(\mu,\nu^{\Theta}_{\omega})$ on $\mathbb{R}^{d}\times\mathbb{R}$ with marginal constraints $\mu$ and $\nu^{\Theta}_{\omega}$, generated by the unprojected data $X$ and the projected generator output $\omega^{T}G_{\Theta}(Z)$, respectively. That is, a $\pi^{\prime}\in\Pi^{\prime}$ satisfies the marginal conditions: $(\text{proj}^{1}_{\pi^{\prime}}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d})\#\pi^{\prime}=\mu$, and $(\text{proj}^{2}_{\pi^{\prime}}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R})\#\pi^{\prime}=\nu^{\Theta}_{\omega}$. The new sliced WGAN problem is
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi^{\prime}\in\Pi^{\prime}(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi^{\prime}}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\right)^{\frac{1}{q}}.$ (7)
Comparing (7) with (6), the optimal transport function for the inner discriminator problem of (7) is different from that of (6) due to the definitions of sets $\Pi^{\prime}$ and $\Pi_{\omega}$. The main difference is on their first marginal condition. For a $\pi_{\omega}\in\Pi_{\omega}$, it satisfies $(\text{proj}^{1}_{\pi_{\omega}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R})\#\pi_{\omega}=\mu_{\omega}$, where $\mu_{\omega}$ is the projected data distribution, while a $\pi^{\prime}\in\Pi^{\prime}$ satisfies $(\text{proj}^{1}_{\pi^{\prime}}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d})\#\pi^{\prime}=\mu$, where $\mu$ is the original data distribution.
Note that for all generator functions, the corresponding optimal values from original (6) and the new (7) with unprojected $\mu$ are non-negative. Thus, we say that a sequence of generator functions is asymptotically optimal for (6) if the following holds: there exists a sequence of generator functions $G_{\Theta}(\cdot)\in\mathbb{R}^{d}$ such that the corresponding values of (6) go to zero as $d\rightarrow\infty$. The same definition applies to (7). In Sec. III-B, we identify the asymptotically optimal generator for both original and unprojected sliced WGAN problems (6) and (7).
TABLE I: Comparison to [12, 29, 3], where $q$ is WGAN order, and $d$ is data dimension
Paper/Theorem Non-linear generator Closed-form solution Continuous data distribution WGAN
[12] X V Gaussian $q=2$
[29] X V Gaussian $q=2$ regularized
[3] V X Gaussian/Uniform $q=1$
Theorem 1 V V No limit $q=2,d=1$
Theorem 2 V V No limit $q=2$ sliced
Theorem 3 V V No limit $q=2$ unprojected sliced